Thomas has an enormous stamp collection. He has wallpapered one wall of his house with stamps that are square with sides that are one and one-quarter inches long. How many stamps were used to cover the wall, which is 10 feet tall and
25 feet long?

Found 2 solutions by Alan3354, mananth:
Answer by Alan3354(69443)

Thomas has an enormous stamp collection. He has wallpapered one wall of his house with stamps that are square with sides that are one and one-quarter inches long. How many stamps were used to cover the wall, which is 10 feet tall and 25 feet long?
---------------
Assuming all the area of the wall is covered:
The area of the wall is 10*25 = 250 sq feet = 250*144 sq inches
Wall area = 36000 sq inches
------------
Stamp area = 1*1.25
# of stamps = 36000/1.25
= 28800 stamps

Answer by mananth(16946)

Number of stamps stuck = area of wall / area of stamp
area of wall = (10*12 )*(15*12) inches
= 21600 square inches.
..
Area of stamp = 1*1.25 square inches
..
21600/1.25
= 17,280 stamps